OpenStudy (anonymous):

someone please help ????? Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y = 1/3x - 8 y = -3x + 12 y = 3x - 24 y = -1/3x - 4

OpenStudy (anonymous):

try to draw manually

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

Calculate the slope of the line perpendicular to y = 3x - 6 Slope of the given line is 3..so, slope of its perpendicular will be -1/3... So, now you have slope -1/3 and point (6,-6) Find the equation of line..can you? use slope-point form of line to get it

OpenStudy (anonymous):

so it would be y = 3x - 24 ?

OpenStudy (amriju):

nooooooooooo

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

well its either that or y = -1/3x - 4

OpenStudy (amriju):

thats going by the options....u have workd it it out..?

OpenStudy (anonymous):

Yes its D By using slope-point form (y+6) = -1/3 (x-6) y+6 = -1/3x + 2 y = -1/3x - 4

OpenStudy (anonymous):

its a big mess of numbers so not really.

OpenStudy (anonymous):

i knew it ! thanks @AbhimanyuPudi

OpenStudy (anonymous):

welcome :-)

OpenStudy (anonymous):

can u help me with other problem ?

OpenStudy (anonymous):

Sure..